Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
July 23, 2007

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Chris Nicolich, Amy Polk, Mark Sullivan, Stephen Williams
Announcements
  1. Cynthia is considering wallpapering her refrigerator after reading an article about it online.
  2. Frank, the yellow-belly water snake that was recently spending time in Bobbie's yard and porch, has been killed.
  3. The Library may be reorganizing - with Carol Turner's impending retirement, the number of interim positions and a new dean - it seems likely. The subject is to be discussed at the management retreat in August.
  4. Supplies - new budget received. Each department is to do their own ordering and tracking of expenses. Debra will have a list in the supply closet where staff can post needed items. Orders will be place on the 15th and 30th of the month, unless there is a critical need for immediacy. The budget is small, so please conserve.
Review Minutes of 7/16/07
  1. Ann #2 - Remove the new (inaccurate and inactive) phone number for Systems.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders are ongoing - nothing over $1000.
Liaisons
  1. Training to start again in August. Amy is looking to nail down the best week.
Networking
  1. Exchange server - it has been a royal pain in the past few weeks. There have been problems with the mobile synching, webmail, DNS and the Domain Controller queues are backed up. Winston notes, in a possibly related Exchange problem, Blake Landor set his "out of office" message, but it is only effective with in house mail. Bobbie suggests that there could be a problem with the back-up client on Exchange - the remote agent. We could take it off and see how it goes. We will be keeping the Exchange server for the foreseeable future. How long is unknown - an upgrade is questionable right now as we have to figure out the current problems.
  2. Verify Patron - ready for another test - can't find SQL or failover servers.
  3. Backup exec problems with SQL servers: NT19 daily failed Friday night. NT 18 and 2K3SQL is giving the failure error, "no disk space on the C: drive".
  4. Xerox - the usual issues.
  5. Logan will be out at VM ware for the SAN four days next week.
  6. Server updates - NT25 on hold.
Software Projects Status
  1. Grover review on schedule for this Friday morning.
  2. Tracking database - one more change forthcoming. Erich has assigned someone doing acceptance testing.
  3. Greenstone - no major changes.
  4. Patron Updates - Winston released changes for v.18 and is currently loading fall as well as summer students.
  5. ILL - changing authentication from barcode to UFID.
Students and Department
  1. Vera's last day is Thursday, July 26. Mindy will pick up more hours for the rest of the summer semester.
  2. Luiza will return in the fall. Fall semester starts 8/23.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page